### v1

@@ B.3 Statements and proofs of claims in Remark 4
For the case where $\beta$ is odd, consider any positive integer $k$, and consider $\beta$ such that $b_{0}^{2k+1}<\beta<b_{1}^{2k+1}$. For any $m$, we will show that the null is locally-dense by considering intervals $x_{1},y_{1}],\cdots,[x_{m},y_{m}]\subset\mathcal{B}$ where
@@
For the case where $\beta$ is even, we will show that for any real number $0<\beta<\max\{b_{0}^{2k},b_{1}^{2k}\}$ and any continuous random variable $X$ with probability density function $p$ supported on $m$ intervals $x_{1},y_{1}],\cdots,[x_{m},y_{m}]\subset\mathcal{B}$ where
@@
We next consider the statement in Proposition 8 replacing $\Omega_{\text{cts}}(\mathbb{R})$ with $\Omega_{\text{cts}}([b_{0},b_{1}])$ for some interval $b_{0},b_{1}]$ and considering any $q\in(b_{0},b_{1})$. Let $X$ be a continuous random variable with probability density function $f$ supported on $m$ intervals $x_{1},y_{1}],\cdots,[x_{m},y_{m}]\subset[b_{0},b_{1}]$ where $\max_{1\leq i\leq m}|y_{i}-x_{i}|<\min\{\frac{q-b_{0}}{m+1},\frac{b_{1}-q}{m+1}\}$. We will show that for any $q\in(0,1)$, we can find an $\alpha\in(0,1]$ and a probability density function $g$ on the real line with the $p$-th quantile equal to $q$ and $g(x)|_{\bigcup_{i=1}^{m}[x_{i},y_{i}]}=\alpha f(x)$.
@@
Now we can choose $\alpha=a>0$ small enough such that $\alpha\operatorname{\mathbb{P}}[X\leq q]<p$. In order for $\operatorname{\mathbb{P}}[X_{\alpha}\leq q]=p$, we need $\operatorname{\mathbb{P}}[Y\leq q]=\frac{p-a\operatorname{\mathbb{P}}[X\leq q]}{1-\alpha}.$ This can be achieved by letting $h(x)=\frac{p-a\operatorname{\mathbb{P}}[X\leq q]}{(1-\alpha)(y-x)}\mathbb{1}_{[x,y]}+(1-\frac{p-a\operatorname{\mathbb{P}}[X\leq q]}{(1-\alpha)})\frac{1}{(z-w)}\mathbb{1}_{[w,z]}$. ∎
Proposition 10.

### v2

@@ B.3 Statements and proofs of claims in Remark 4
For the case where the moment order is odd, consider any positive integer $k$, and consider $\beta$ such that $b_{0}^{2k+1}<\beta<b_{1}^{2k+1}$. For any $m$, we will show that the null is locally-dense by considering intervals $x_{1},y_{1}],\cdots,[x_{m},y_{m}]\subset\mathcal{B}$ where
@@
For the case where the moment order is even, we will show that for any real number $0<\beta<\max\{b_{0}^{2k},b_{1}^{2k}\}$ and any continuous random variable $X$ with probability density function $p$ supported on $m$ intervals $x_{1},y_{1}],\cdots,[x_{m},y_{m}]\subset\mathcal{B}$ where
@@
We next consider the statement in Proposition 8 replacing $\Omega_{\text{cts}}(\mathbb{R})$ with $\Omega_{\text{cts}}([b_{0},b_{1}])$ for some interval $b_{0},b_{1}]$ and considering any $q\in(b_{0},b_{1})$. Let $X$ be a continuous random variable with probability density function $f$ supported on $m$ intervals $x_{1},y_{1}],\cdots,[x_{m},y_{m}]\subset[b_{0},b_{1}]$ where $\max_{1\leq i\leq m}|y_{i}-x_{i}|<\min\{\frac{q-b_{0}}{m+1},\frac{b_{1}-q}{m+1}\}$. We will show that for any $p\in(0,1)$, there exist $\alpha\in(0,1]$ and a probability density function $g$ on $b_{0},b_{1}]$ such that $q$ is the $p$-th quantile of $g$ and $g(x)|_{\bigcup_{i=1}^{m}[x_{i},y_{i}]}=\alpha f(x)$.
@@
Now we can choose $\alpha>0$ small enough such that $\alpha\operatorname{\mathbb{P}}[X\leq q]<p$. In order for $\operatorname{\mathbb{P}}[X_{\alpha}\leq q]=p$, we need $\operatorname{\mathbb{P}}[Y\leq q]=\frac{p-\alpha\operatorname{\mathbb{P}}[X\leq q]}{1-\alpha}.$ This can be achieved by letting $h(x)=\frac{p-\alpha\operatorname{\mathbb{P}}[X\leq q]}{(1-\alpha)(y-x)}\mathbb{1}_{[x,y]}+(1-\frac{p-\alpha\operatorname{\mathbb{P}}[X\leq q]}{(1-\alpha)})\frac{1}{(z-w)}\mathbb{1}_{[w,z]}$. ∎
Proposition B.10.